### v1

# MROVSeg: Breaking the Resolution Curse of Vision-Language Models in Open-Vocabulary Semantic Segmentation

Abstract
Open-vocabulary semantic segmentation aims to segment and recognize semantically meaningful regions based on text-based descriptions during inference. A typical solution to address this task is to leverage powerful vision-language models (VLMs), such as CLIP, to bridge the gap between open- and close-vocabulary recognition. As VLMs are usually pretrained with low-resolution images (e.g. $224\times 224$), most previous methods operate only on downscaled images. We question this design as low resolution features often fail to preserve fine details. Although employing additional image backbones for high-resolution inputs can mitigate this issue, it may also introduce significant computation overhead. Therefore, we propose MROVSeg, a multi-resolution training framework for open-vocabulary semantic segmentation with a single pretrained CLIP backbone, that uses sliding windows to slice the high-resolution input into uniform patches, each matching the input size of the well-trained image encoder. Its key components include a Multi-Res Adapter, which restores the spatial geometry and grasps local-global correspondences across patches by learnable convolutional and scale attention layers. To achieve accurate segmentation, we introduce Multi-grained Masked Attention scheme to aggregate multi-grained semantics by performing cross-attention between object queries and multi-resolution CLIP features within the region of interests. Through comprehensive experiments, we demonstrate the superiority of MROVSeg on well-established open-vocabulary semantic segmentation benchmarks, particularly for high-resolution inputs, establishing new standards for open-vocabulary semantic segmentation.
1 Introduction
(a) Previous open-vocabulary semantic segmentation.
(b) MROVSeg open-vocabulary semantic segmentation.
Figure 1: Comparison between other training frameworks and MROVSeg. Previous methods (a) adopt additional image backbone to provide mask feature. The mask prediction is class-unaware. Our method (b) provide multi-resolution CLIP feature for both mask decoding and mask classification, and the whole framework is class-aware.
Semantic segmentation is a fundamental computer vision problem which seek to group image pixels into semantically meaningful segments and recognize their categories. Previous research in semantic segmentation [25, 6, 9] mainly focus on segmenting objects from a pre-defined semantic categories set based on the dataset. These methods typically rely on human-annotated masks and class labels, which is costly and unscalable in terms to the number of categories. To this end, the task of open-vocabulary semantic segmentation has been introduced [39, 16, 21], which aims to segment semantic pixels belonging to arbitrary classes beyond pre-defined categories and dataset.
In recent years, large-scale vision-language pretrained models (VLMs), such as CLIP [28] and ALIGN [19], have demonstrated remarkable generalization capabilities for recognizing open-vocabulary categories. This motivate the research community to investigate the potential of VLMs in open-vocabulary semantic segmentation. To address the discrepancy between the per-pixel semantic requirements and the image-level labels provided by VLMs, initial studies [11, 44, 45] modified CLIP model by removing its final pooling layer to obtain dense category embeddings for per-pixel classification. However, these approaches typically necessitate fine-tuning VLMs on a base segmentation dataset with limited images and categories, which is demonstrated [44] to impair the transferability of VLM features, leading to unsatisfactory zero-shot performance on downstream tasks.
Recent approaches [35, 14, 36, 30, 24] reformulate open-vocabulary semantic segmentation as a region-level recognition problem. These methods typically adopt two branch meta architecture: one branch extract image feature and generate mask proposals, and the other branch classifies the predicted proposals with pretrained VLM. Although these methods are promising, we note their following limitations. First, due to pretrained VLMs exhibit inferior size adaptability, most of open-vocabulary semantic segmentation methods (e.g. [14, 36, 22, 35, 34, 8, 30]) so far need to downsample images to fit the pretrained resolution (e.g. $224\times 224$) of VLM to perform region-level recognition (as in Fig. 1 (a)). However, low-resolution input usually lacks segmentation details. Although naïvely applying sliding window inference [11, 35] could partly compensate for the details, the spatial structure across windows is corrupted and the local-global modeling is also absent.
In light of the limitations and challenges faced by previous methods, we propose MROVSeg, a VLM-based Multi-Resolution training framework for Open-Vocabulary Semantic Segmentation. As illustrated in Fig.  1 (b), first, MROVSeg uses downsampled low-resolution images as VLM input to extract global low-resolution feature. Second, MROVSeg split the high-resolution images into slices and feeds them to VLM to extract detailed high-resolution features. The key components of MROVSeg contain a Multi-Res Adapter, in which we employs depthwise convolution layers to restore the spatial geometry across slices. To effectively capture global long-range context, inspired by previous multi-scale training framework [5, 18, 37], we employ a image-dependent Scale-aware Attention [5] to dynamically adjust the fusion weight of high-resolution and low-resolution VLM features based on their relevance and trustworthiness. The resulting multi-resolution features are fused hierarchically then employed for precise mask proposals generation.
To achieve accurate mask class recognition, we propose a Multi-grained Masked Attention mechanism. The core insight is to reuse the CLIP [CLS] token, and manipulate its attention map on multi-resolution features in CLIP attention layers with resolution-aware attention masks. We find this resolution-aware design can enforce the low- and high-resolution attention map focus on global contexts and spatial details respectively and thus effectively aggregate multi-grained semantics.
With extensive experiments on well-established open-vocabulary semantic segmentation benchmarks, such as Pascal Context-59, Pascal Context-459 [26], ADE-150, ADE-847 [43], Pascal VOC [15], we are delighted to report that our method surpasses current state-of-the-art methods by a significant margin on standard benchmarks (+2.4 mIoU%), demonstrating the advancements of MROVSeg in the domain of open-vocabulary semantic segmentation. Our contributions can be summarized as follows:
  • We propose a novel end-to-end multi-resolution training framework to tackle the task of open-vocabulary semantic segmentation. It enables improved open-vocabulary segmentation by leveraging multi-resolution vision-language features.
  • A multi-grained masked attention scheme is proposed to effectively aggregate regional and universal semantics on multi-resolution vision-language features.
  • The efficacy of our method is confirmed by achieving state-of-the-art performance by evaluating our proposed approach on five well-established open-vocabulary semantic segmentation benchmarks.
2 Related work
Figure 2: The overall pipeline of MROVSeg. For an high-resolution input image, its downsampled image and are fed into CLIP visual encoder to extract multi-resolution CLIP features. The multi-resolution CLIP features are injected into Multi-Res Adapter to produce multi-scale features for mask decoder.
Pretrained Vision-language Models. Recently, large-scale pretrained, contrastive learning based methods [28, 19] demonstrate powerful open-vocabulary image classification capability by learning shared image-text feature representations. Pretrained CLIP [28] has been generalized to many downstream computer vision tasks such as object detection [42], image caption [17], image generation [27] and image segmentation [36, 2, 34, 38, 35]. Our method invokes natural language perceptual ability of pretrained VLMs, aiming to explore their application boundary in open vocabulary semantic segmentation tasks.
Semantic Segmentation. Classic closed-vocabulary semantic segmentation interprets category concepts in a predefined label sets and generally performs segmentation on a specific dataset. In the past decade, FCN [25] and its variants [40, 6, 7, 41, 33] have dominated the field for several years by using per-pixel classification framework. Recently, MaskFormer [9, 10] explore to reformulate the semantic segmentation as a mask classification problem. Our approach is the extension of the traditional semantic segmentation methods, enabling them to segment a much broader range of semantic categories and adapt to real-world scenarios.
Open Vocabulary Image Segmentation. Open vocabulary image segmentation aims at segmenting and recognizing semantic regions described by text. Pioneering works [2, 31] use learned language embedding to align the feature space of class texts and visual semantics. Recently, some works [35, 14, 22, 24] develop two-stage training approaches, i.e. first train a stand-alone model to generate class-agnostic masks, and then adapt CLIP models to classify the mask proposals. More recently, some end-to-end frameworks emerge in the community, which unify mask generation and region classification into the same model. SAN [36] propose a lightweight side adapter to generate mask proposals and attention biases that act on CLIP attention layers. ODISE [34] employs a stable diffusion UNet [29] to generate mask proposals and adopt frozen CLIP model to enhance the open-vocabulary recognition capability. With the extracted dense pixel semantic embedding [44], CAT-Seg [11] proposes to finetune CLIP with cost aggregation. EBSeg [30] integrate powerful Segment Anything Model [20] into CLIP-based framework with image embedding balancing.
Discussion with Previous Methods. Our method MROVSeg is inspired by [36, 38], but has significant differences: (1) MROVSeg, which adapts multi-resolution CLIP feature to assist semantic segmentation by Multi-Res Adapter, is a multi-resolution training framework, and is different from all previous open-vocabulary semantic segmentation methods. (2) The introduced Muti-grained Masked Attention scheme explicitly enforces the mask class recognition to aggregate both local and global semantics, which takes advantage of the internal consistency between multi-resolution features.
3 Method
3.1 Problem Definition
In open vocabulary semantic segmentation task setting [35, 22], an image $I\in\mathbb{R}^{3\times H\times W}$ is operated by a segmentation model $\mathcal{P}_{\theta}$ with parameter $\theta$ to produce a set of masks associated with $K$ semantic categories:
$\{m_{i},y_{i}\}_{i=1}^{K}=\mathcal{P}_{\theta}(I).$ (1)
The segmentation model $\mathcal{P}_{\theta}$ is trained on a base segmentation dataset (e.g., COCO [3]) annotated with a fixed label set $\mathcal{Y}_{\mathrm{train}}=\{y_{i}\}_{i=1}^{K_{\mathrm{train}}}$ of $K_{\mathrm{train}}$ semantic categories. And during test, model $\mathcal{P}_{\theta}$ is expected to segment objects of category set $\mathcal{Y}_{\mathrm{test}}$, which generally contains novel categories ($\mathcal{Y}_{\mathrm{test}}\not\subset\mathcal{Y}_{\mathrm{train}}$).
3.2 Motivation
Accurate semantic segmentation needs high-resolution image inputs. Due to low-resolution images used in vision-language pretraining, previous open-vocabulary methods employ extra image backbone (such as SAM [30] and ResNet [24]) to provide segmentation detail. Although recent studies [32, 38] adapt convolution-based CLIP model for high-resolution training, but directly apply these methods to ViT-based CLIP model results in suboptimal performance [36] due to undesirable size adaptability of ViT. To this end, we propose MROVSeg, a ViT-based training framework to provide multi-resolution vision-language features for open-vocabulary semantic segmentation.
3.3 Framework Overview
The overview of our training framework MROVSeg, for open-vocabulary semantic segmentation is shown in Fig. 2. At a high-level, an image is downsampled and padded to a low-resolution (such as $224\times 224$) and processed by a pretrained CLIP ViT to extract global feature. To capture high-resolution local details, the high-resolution image is split into slices and input to the shared CLIP ViT encoder. These multi-resolution features are then concatenated with learnable queries and fed into a Multi-Res Adapter(Sec. 3.4), to produce the fused features, query features, and attention masks used for mask classification.
Figure 3: Multi-Res Adapter. The slice features from CLIP layer 0 $\{\mathbf{P}_{i}^{0}\}_{i=1}^{S}$ are concatenated with learnable queries and fed to ViT Blocks. The slice features from various CLIP layers are first adapted by MRF module to restore spatial geometry and capture long-range global contexts, then are injected to the intermediate ViT Blocks. The final output visual tokens and projected queries are utilized for downstream mask prediction and classification.
3.4 Multi-Res Adapter
As depicted in Fig. 3, the Multi-Res Adapter adapts multi-resolution CLIP features for segmentation task. Denote the slice (high resolution) features of $l$-layer as $\{\mathbf{P}_{i}^{l}\}_{i=1}^{S}$, the global feature (low resolution) as $\bar{\mathbf{P}}_{i}^{l}$, where $\mathbf{P}_{i}\in\mathbb{R}^{L\times D}$, $S$ is the slice number, $L=H\times W$ is the token length, and $D$ is the channel number. Naïvely concatenating the high resolution slice features for the subsequent segmentation prediction is promising, but have two defects: 1) the spatial geometry, such as the positional information, is corrupted across slice features $\{\mathbf{P}_{i}^{l}\}_{i=1}^{S}$; 2) the long-range dependencies and global context is missing. Thus, we propose Multi-Res Adapter to effectively restore spatial geometry of slice features and capture long-range context from global feature. In Multi-Res Adapter, the 00-th slice features $\{\mathbf{P}_{i}^{0}\}_{i=1}^{S}$ are firstly concatenated with learnable queries $\mathbf{Q}\in\mathbb{R}^{N\times D}$ and input to the vanilla ViT blocks to build the query features for each objects. Then for target fusion layer $l$, the slice features $\{\mathbf{P}_{i}^{l}\}_{i=1}^{S}$ and global feature $\bar{\mathbf{P}}_{i}^{l}$ are fused through a Multi-Res Fusion (MRF) Module and then injected into the ViT branch.
Multi-Res Fusion (MRF) Module first reshape the global feature to $\bar{\mathbf{H}}^{l}\in\mathbb{R}^{H\times W\times D}$, and restore slice features to $\mathbf{H}^{l}\in\mathbb{R}^{m*H\times n*W\times D}$, where $S=m\times n$. To retain the spatial geometry of high-resolution features, we employ depth-wise separable convolutions to fuse the restored feature. To effectively model the local-global correspondence, we train a Scale-aware Attention [5] to fuse multi-res features into $\mathbf{F}^{l}\in\mathbb{R}^{H\times W\times D}$ as the fused feature
$\mathbf{F}^{l}=\texttt{Up}(a_{l})\odot\texttt{DConv}(\mathbf{H}^{l})+\texttt{Up}((1-a_{l})\odot\bar{\mathbf{H}}^{l}).$ (2)
Then $\mathbf{F}^{l}$ is added to the visual tokens in Multi-Res Adapter. The scale attention decoder $f_{A}^{l}$ learns to predict the scale attention $a_{l}=\sigma(f_{A}^{l}(\mathbf{H}_{l}))\in[0,1]^{H\times W\times D}$ for layer $l$ to weigh the trustworthiness of low resolution context and high resolution detail. The sigmoid function $\sigma$ ensures $a_{l}$ weight in $[0,1]$, where $1$ means $a_{l}$ focus on high resolution detail. In practice, we empirically select the features from a CLIP layer set $\{l_{j}\}_{j=1}^{N_{j}}$ to apply in Multi-Res Adapter. For instance, for the model based on CLIP ViT-L model, $l\in\{\texttt{stem},6,12,18\}$. The fused features $\{\textbf{F}^{l_{j}}\}_{j=1}^{N_{j}}$ are used for hierarchical mask decoding. The final layer output slice features $\{\mathbf{P}_{i}^{0}\}_{i=1}^{S}$ are restored to $\mathbf{H}\in\mathbb{R}^{m*H\times n*W\times D}$ as the visual feature for hierarchical mask decoding and multi-grained masked attention. And the output queries $\mathbf{Q}$ are projected as the query feature
$\mathbf{Q}_{f}=\texttt{MLP}^{\mathrm{Q}}(\mathbf{Q}),$ (3)
for hierarchical mask decoding and multi-grained masked attention.
3.5 Mask Prediction
[…]
$\mathrm{\mathbf{H}}_{\texttt{pix}}=\texttt{MLP}^{\texttt{pix}}(\mathrm{\mathbf{H}}_{\texttt{up}}),$ (4)
$\mathrm{\mathbf{M}}_{\texttt{mask}}=\mathrm{\mathbf{Q}}_{f}\times\mathrm{\mathbf{H}}_{\texttt{pix}},$ (5)
where the query feature $\mathrm{\mathbf{Q}}_{f}$ is from the Multi-Res Adapter which described in Sec. 4.3. $\mathrm{\mathbf{M}}_{\texttt{mask}}\in[0,1]^{N\times H\times W}$ is the mask prediction, and we omit the sigmoid function in Eq.5.
3.6 Mask Classification
Decoupled Attention Mask Decoding. CLIP models, like general Vision Transformers, perform open-vocabulary image-level recognition through [CLS] token. Our goal is to recognize the various predicted mask proposals without changing any parameter of CLIP by guiding the attention maps of [CLS] token on the region of interests. To achieve this goal, we follow MaskCLIP [14] and SAN [36] to reuse the original CLIP [CLS] token in the intermediate layer. To sufficiently aggregate multi-grained semantics from CLIP, we first duplicate the [CLS] token to query number $N$ and create learnable positional embedding for them, dubbed as the $\mathrm{\mathbf{X}_{\texttt{prop}}}\in\mathbb{R}^{N\times D}$. We aim to enforce $\mathrm{\mathbf{X}_{\texttt{prop}}}$ to extract the global and local semantics from low- and high-resolution CLIP feature respectively. Thus, for visual feature $\mathbf{H}$, we first extract global contexts with max pooling $\bar{\mathbf{H}}=\texttt{MaxPool}(\mathbf{H})$ and train MLPs project them to attention space
$\mathbf{A}_{\mathrm{local}}=\texttt{MLP}^{\texttt{L}}(\mathbf{H}),\quad\mathbf{A}_{\mathrm{global}}=\texttt{MLP}^{\texttt{G}}(\bar{\mathbf{H}}),$ (6)
where $\mathbf{A}_{\texttt{global}}\in\mathbb{R}^{H\times W\times D^{\prime}},\mathbf{A}_{\texttt{local}}\in\mathbb{R}^{m*H\times n*W\times D^{\prime}}$ denote the local and global attention features respectively. Then we decode local and global per-head attention masks by the inner product with
$\mathbf{M}_{\mathrm{local}}=\mathbf{Q}_{f}\times\mathbf{A}_{\mathrm{local}}^{\mathrm{T}},\enspace\mathbf{M}_{\mathrm{global}}=\mathbf{Q}_{f}\times\mathbf{A}_{\mathrm{global}}^{\mathrm{T}},$ (7)
where $\mathbf{Q}_{f}$ is the output query feature described in Sec.3.4. We show this decoupled resolution-aware attention decoding benefit the multi-grained aggregation in Fig.7.
Multi-grained Masked Attention. As shown in Fig.4, we perform cross attention to update the $\mathrm{\mathbf{X}}_{\texttt{prop}}$ with multi-resolution CLIP features, with the predicted attention masks $\mathbf{M}_{\mathrm{local}}$ and $\mathbf{M}_{\mathrm{global}}$,
$\!\!\mathbf{X}^{l+1}_{\texttt{prop}}\!=\!\texttt{softmax}(\mathbf{Q}^{l}_{\texttt{prop}}\!\!\left[\begin{array}[]{c}\mathbf{K}^{l}_{\texttt{LR}}\\ \mathbf{K}^{l}_{\texttt{HR}}\\ \end{array}\right]^{\!\mathrm{T}}\!\!\!\!+\!\left[\begin{array}[]{c}\mathbf{M}_{\mathrm{global}}\\ \mathbf{M}_{\mathrm{local}}\end{array}\right]^{\!\mathrm{T}}\!\!\!)\!\left[\begin{array}[]{c}\mathbf{V}^{l}_{\texttt{{LR}}}\\ \mathbf{V}^{l}_{\texttt{{HR}}}\end{array}\right]\!+\mathbf{X}^{l}_{\texttt{prop}},$ (8)
where $\mathbf{Q}_{\texttt{prop}}=\mathbf{W}_{\mathrm{q}}^{\mathrm{T}}\mathbf{X}_{\texttt{prop}}$ is query embeddings. Denote the low- and high-resolution CLIP tokens as $\mathbf{X}_{\texttt{LR}}$ and $\mathbf{X}_{\texttt{HR}}$. $\mathbf{K}_{\texttt{LR}}=\mathbf{W}_{\mathrm{k}}^{\mathrm{T}}\mathbf{X}_{\texttt{LR}}$ and $\mathbf{K}_{\texttt{HR}}=\mathbf{W}_{\mathrm{k}}^{\mathrm{T}}\mathbf{X}_{\texttt{HR}}$ are the key embeddings of low- and high-resolution CLIP visual tokens respectively. $\mathbf{V}_{\texttt{LR}}=\mathbf{W}_{\mathrm{v}}^{\mathrm{T}}\mathbf{X}_{\texttt{LR}}$ and $\mathbf{V}_{\texttt{HR}}=\mathbf{W}_{\mathrm{v}}^{\mathrm{T}}\mathbf{X}_{\texttt{HR}}$ are value embeddings. $\mathbf{W}_{\mathrm{q}}$,$\mathbf{W}_{\mathrm{k}}$ and $\mathbf{W}_{\mathrm{v}}$ are projection weights of cross-attention layer. The final output proposal logits $\mathbf{X}_{\texttt{prop}}$ are projected to the shared vision-language space and compute cosine similarity with text embeddings to obtain proposal logits $\mathbf{C}\in\mathbb{R}^{N\times K}$: $\mathbf{C}=\mathbf{X}_{\texttt{prop}}\mathbf{W}_{\mathrm{visual}}\mathbf{W}_{\mathrm{text}}^{\mathrm{T}}\mathbf{X}_{\texttt{text}}^{\mathrm{T}}$, where $K$ is the number of categories, and $\mathbf{W}_{\mathrm{visual}}$ and $\mathbf{W}_{\mathrm{text}}$ are projection weights. Finally, the final segmentation map $\mathbf{S}\in\mathbb{Z}_{+}^{K\times H\times W}$ is produced by
$\mathbf{S}=\mathbf{C}\times\mathbf{M}_{\texttt{mask}}^{\mathrm{T}}.$ (9)
Figure 4: Multi-grained Masked Attention.
4 Experiments
Table 1: Performance comparison with state-of-the-art methods. †: We cite the reproduction performance of these methods[35, 13] trained with full COCO Stuff dataset from previous works[36, 11]. PC: Pascal Context, VOC: Pascal VOC.
Method VL-Model Training Dataset Extra Backbone ADE-847 PC-459 ADE-150 PC-59 VOC-20
Zegformer† [13] CLIP ViT-B/16 COCO Stuff ResNet-101 5.6 10.4 18.0 45.5 89.5
ZSSeg† [35] CLIP ViT-B/16 COCO Stuff ResNet-101 6.9 9.7 21.1 51.9 91.8
OvSeg [22] CLIP ViT-B/16 COCO Stuff ResNet-101c 7.1 11.0 24.8 53.3 92.6
SAN [36] CLIP ViT-B/16 COCO Stuff - 10.7 13.7 28.9 55.4 94.6
CAT-Seg [11] CLIP ViT-B/16 COCO Stuff ResNet-101 8.4 16.6 27.2 57.5 93.7
EBSeg [30] CLIP ViT-B/16 COCO Stuff SAM-B 11.1 17.3 30.0 56.7 94.6
SCAN [24] CLIP ViT-B/16 COCO Stuff ResNet-101 10.8 13.2 30.8 58.4 97.0
SED [32] CLIP ConvNeXt-B COCO Stuff - 11.2 18.6 31.6 57.3 94.4
MROVSeg CLIP ViT-B/16 COCO Stuff - 12.1 19.6 32.0 58.5 95.5
OvSeg [22] CLIP ViT-L/14 COCO Stuff Swin-B 9.0 12.4 29.6 55.7 94.5
MaskCLIP [14] CLIP ViT-L/14 COCO Panoptic ResNet-50 8.2 10.0 23.7 45.9 -
ZSSeg† [35] CLIP ViT-L/14 COCO Stuff ResNet-101 7.1 10.2 21.7 52.2 92.3
SAN [36] CLIP ViT-L/14 COCO Stuff - 13.7 17.1 33.3 60.2 95.5
ODISE [34] CLIP ViT-L/14 COCO Panoptic Stable Diffusion 11.0 13.8 28.7 55.3 -
CAT-Seg [11] CLIP ViT-L/14 COCO Stuff Swin-B 10.8 20.4 31.5 62.0 96.6
EBSeg [30] CLIP ViT-L/14 COCO Stuff SAM-B 13.7 21.0 32.8 60.2 96.4
SCAN [24] CLIP ViT-L/14 COCO Stuff ResNet-101 14.0 16.7 33.5 59.3 97.2
FC-CLIP [38] CLIP ConvNeXt-L COCO Panoptic - 14.8 18.2 34.1 58.4 95.4
SED [32] CLIP ConvNeXt-L COCO Stuff - 13.7 22.1 35.3 60.9 96.1
MROVSeg CLIP ViT-L/14 COCO Stuff - 16.1 24.1 36.9 64.1 97.6
4.1 Setting
We train our models on COCO-Stuff [3] dataset which comprises 164K images with densely annotated masks spanning 171 categories. We first evaluate MROVSeg on five well-established mainstream open-vocabulary semantic segmentation benchmarks for standard evaluation. We further evaluate MROVSeg on Cityscapes [12] benchmarks to explore the ability of handling high-resolution image input. We follow common practice [35, 2] to measure the segmentation performance by mean intersection over union (mIoU) score.
Datasets. The standard benchmarks contains three dataset: ADE [43], Pascal Context [26], and Pascal VOC [15]. The ADE dataset contains around 20K and 2K images for training and validation, respectively. This dataset is annotated with 150 and 847 categories, resulting in two separate segmentation benchmarks, namely ADE-150 and ADE-847. Similarly, the Pascal Context dataset has 5K images for both training and validation. It is annotated with 59 and 459 classes, forming two benchmarks known as PC-59 and PC-459. The Pascal VOC dataset comprises 1464 and 1449 images for training and validation, encompassing annotated masks across 20 semantic categories.
Implementation Details. We adopt the vanilla ViT block as transformer block in Multi-Res Adapter, and we use 6 blocks with 12 attention heads, 100 query tokens, feature dimension is 768 as default. We choose OpenAI pretrained ViT-based CLIP [28] models in all experiments for better reproducibility. We empirically choose [CLS] token from CLIP layer 9, CLIP ViT-B/16 model, and subsequent layers for Multi-grained Masked Attention. i.e., we use last 3 blocks for Multi-grained Masked Attention. For CLIP ViT-L/14 model, we [CLS] token from CLIP layer 18, and use subsequent layers for Multi-grained Masked Attention. Our models are trained with image input resolution $640\times 640$ , and slice and downsample to $320\times 320$ to fit the CLIP input resolution. More implementation details are in Appendix Sec. B.
Figure 5: Visualization of image resolution distribution histogram of the datasets in Tab.1.
Figure 6: Effect of scaling up backbone on Cityscapes.
Figure 7: Effect of decoupled attention decoding for multi-grained semantics. With single attention mask decoding, the model is unable to recognize the class painting (h) in the wall even with proper predicted proposals (g) since the spatial cues are overwhelmed by background noise (i). Our decoupled attention mask decoding effectively split the global and local semantics in LR (c) and HR (f) attention masks, performing enhanced recognition (e).
4.2 Standard Evaluation
Comparisons with State-of-the-Art Methods. In Tab. 1, we compare MROVSeg with current state-of-the-art open-vocabulary semantic segmentation methods on standard benchmarks. First of all, our method significantly outperforms other state-of-the-art methods all various open-vocabulary semantic segmentation benchmarks. Specifically, our method supasses state-of-the-art method SAN [30] with the same CLIP backbones on all benchmarks by remarkable margins ($+1.4\%$ mIoU for ADE-847, $+5.9\%$ mIoU for PC-459, $+3.1\%$ mIoU for ADE-150, $+3.1\%$ mIoU for PC-59, and $+0.9$ for VOC-20 with CLIP ViT-B backbone, $+2.4\%$ mIoU for ADE-847, $+7.0\%$ mIoU for PC-459, $+3.6\%$ mIoU for ADE-150, $+3.9\%$ mIoU for PC-59, and $+2.1$ for VOC-20 with CLIP ViT-B backbone). Compared to methods with additional image backbones, our model outperforms EBSeg (with SAM) by $+0.9,+1.0,+3.2,+1.2,+1.1$ mIoUfor ADE-847, PC-459, ADE-150, PC-59, and VOC-20 respectively. In addition, our models outperforms other convolution-based high-resolution trained method FC-CLIP and SED. Fig. 8 shows the qualitative comparison between MROVSeg and state-of-the-art methods (SAN [36] and EBSeg [30]). Evidently, MROVSeg can segment objects more precisely (the first row, class sofa), and provide more detailed mask prediction (the second and third row, more accurate object boundaries).
Capability of Handling High-Rsolution Images. Fig.6 show the image size distribution of the evaluated in Tab.1. It is noteworthy that the size distribution of the test datasets are highly overlapped with COCO, where the image size primarily distributed from $[250,800]$. To evaluate the ability of handling high-resolution image input, we train MROVSeg on COCO-Stuff with $1024\times 1024$ image resolution and compare the performance scalability with convolution-based high-resolution pretrained method FC-CLIP on Cityscapes benchmark. Notably, with high resolution inputs ($2048\times 1024$), MROVSeg reaches comparable performance while scaling up the backbone, which greatly enhanced than other ViT-based methods (EBSeg and SAN). The size adaptability of ViTs have been demonstrated to be worse than ConvNets [38]. Conversely, with our specifically designed multi-resolution training, MROVSeg is able to match the performance of convolution-based method on high-resolution inputs.
Table 2: Ablation study of components. We show the effects of integrating each modules into the baseline.
Multi.Ada. Mul.Ma.Att. Hier.Dec. PC-59 A-150 A-847 PC-459 VOC-20
52.3 25.4 9.0 10.8 93.6
55.6 28.7 10.5 14.9 94.7
57.1 30.5 11.3 18.0 95.1
58.5 32.0 12.1 19.6 95.5
Table 3: Effect of different crop ratio $p$. We also report #GFLOPS and GPU memory consumption (MB). For $p>0.5$, we adopt overlapped slicing.
$p$ Mem. GFLOPS A-150 PC-59 A-847 PC-459 VOC-20
00 6768 116.5 27.6 53.7 10.1 15.7 94.7
$1.0$ 13493 293.1 27.8 54.5 9.9 15.0 94.4
$0.75$ 16261 225.7 30.5 54.3 10.2 15.9 95.1
$0.625$ 12942 184.5 32.1 57.9 12.0 19.3 95.7
$0.5$ 9779 142.8 32.0 58.5 12.1 19.6 95.5
$0.25$ 7560 92.5 28.5 56.2 10.5 17.0 94.7
Table 4: Ablation study on various designs in Multi-Res Adapter. Multi-Res Adapter contains Multi-Res Fusion (MRF) module and ViT blocks. We first show the impact of spatial fusion method (a) and multi-resolution fusion (b). Then we give the impact of number of blocks (c), number of fusion CLIP layer (ViT-B model) (d), and the channel width (e). Finally, we show the impact of query number (f). The default setting are marked underline.
(a) Spatial Fusion A-847 PC-459 A-150 PC-59 VOC-20
Concat. 11.8 18.0 31.5 57.0 94.6
Depth.Conv. 12.1 19.6 32.0 58.5 95.5
(b) Multi-Res Fusion A-847 PC-459 A-150 PC-59 VOC-20
Add 11.7 18.4 31.6 56.9 94.5
Concat. 11.4 18.3 31.0 56.1 94.8
Scale.Attn. 12.1 19.6 32.0 58.5 95.5
(c) Block Num A-847 PC-459 A-150 PC-59 VOC-20
3 8.2 11.5 26.7 53.5 93.8
6 12.1 19.6 32.0 58.5 95.5
9 11.9 19.6 31.5 59.2 95.4
(d) Fusion Layer A-847 PC-459 A-150 PC-59 VOC-20
{3,6,9} 11.2 17.4 31.6 56.5 95.6
{3,6,9,12} 11.7 18.8 31.4 57.0 94.4
{stem,3,6,9} 11.8 19.4 32.1 58.1 94.9
{stem,3,6,9,12} 12.1 19.6 32.0 58.5 95.5
(e) Channel Width A-847 PC-459 A-150 PC-59 VOC-20
384 9.6 16.1 28.4 56.8 94.2
768 12.1 19.6 32.0 58.5 95.5
1024 12.0 19.1 32.4 58.0 95.7
(f) Query Num A-847 PC-459 A-150 PC-59 VOC-20
100 12.1 19.6 32.0 58.5 95.5
200 11.9 18.3 31.8 57.0 93.8
300 12.2 19.3 32.4 57.5 95.2
Figure 8: Qualitative comparison with SAN [36] and EBSeg [30].
4.3 Ablation Study
For all ablation experiments, we employ our CLIP ViT-B/16 based model as our ablated baseline.
Components Analysis. Tab. 2 shows the performance effect of the main components in MROVSeg. The baseline adopts single resolution CLIP plain ViT feature for mask decoding, and use the masked-attention for mask class recognition as in MaskCLIP and SAN. The baseline obtains the mIoU scores of 52.3%, 22.4%, 8.5%, 10.8% and 93.6% for PC-59, ADE-150, ADE-847, PC-459 and VOC-20. When we use multi-resolution images with High-Res Adapter, it has the mIoU scores of 55.6%, 28.7%, 10.5%, 16.8% and 94.7% for PC-59, ADE-150, ADE-847, PC-459 and VOC-20, which outperforms baseline by 3.3%, 6.3%, 2.0%, 6.0% and 1.1%. After use hierarchical features to decode mask, it obtains the mIoU scores of 57.1%, 30.5%, 11.3%, 14.9% and 95.1% for PC-59, ADE-150, ADE-847, PC-459 and VOC-20, which outperforms baseline by 4.8%, 5.1%, 2.3%, 7.2% and 1.5%. Finally we integrate the Multi-grained Masked-Attention to the model, it reaches 58.5%, 32.0%, 12.1%, 19.6% and 95.5% for PC-59, ADE-150, ADE-847, PC-459 and VOC-20, which outperforms baseline by 6.2%, 6.6%, 3.1%, 8.8% and 1.9%.
Single Resolution vs. Multi Resolution. Denote the slice ratio as $p=\frac{slice\;\,resolution}{input\;\,resolution}$, we examine the effect of $p$ value in Tab. 3. Notably, only using single resolution feature, i.e., $p=0$, or directly adopting high-resolution image as CLIP input, i.e., $p=1.0$ lead to significant performance degradation. While some $p$ values with overlapped slicing (such as $0.625$) obtains better performance on some datasets(such as ADE-150 and VOC-20), we choose the default $p$ value as $0.5$ considering computation overhead.
Micro-design Choice of Multi-Res Adapter. We conduct experiment to examine the different micro-design choice in Multi-Res Adapter in Tab. 3. We first perform experiments on MRF module. In (a), we examine different spatial restore strategies. The depthwise convolution is significantly better than concatenation. And we present the impact of local-global modeling methods (b), and the best results on all benchmarks are achieved by scale attention. Then for the ViT block in Multi-Res Adapter, we present the effect of block number in (c), we find that although increase the block number to 9 can bring slight improvement on some benchmarks (PC-59 and PC-459), it would incur more computation cost. In (d), we examine the effect of CLIP layers whose feature we adopt. In (e), we present the impact of channel width we use. Finally, we examine the effect of query number in (f).
Table 5: Efficiency comparisons. We report the and inference FPS of MROVSeg running on a RTX 3090. Training time is mesured with 2 NVIDIA H100.
Method Params. #GFLOPS Inference FPS Training Time
ZSSeg 530.8 22302.1 0.3 15h59min
OVSeg 532.6 19345.6 0.4 17h54min
CAT-Seg 433.7 2121.1 2.0 7h41min
FC-CLIP 221.3 680.0 2.3 2d5h
EBSeg 210.9 867.5 4.7 18h33min
MROVSeg 162.1 640.4 10.5 9h40min
Computation Overhead. We compare the computation overhead of MROVSeg with recent methods [14, 22, 11, 30, 38] in Tab. 5. We measure the number of parameter, giga floating-point operations per second, inference FPS and training time. Our method exhibits strong efficiency among these methods in terms of both training and inference. This is achieved by 1) single CLIP backbone do not need additional image backbones; 2) slice then input strategy avoid quadratic computation cost with regard to input image size.
5 Conclusion
In this paper, we introduce MROVSeg, a multi-resolution training framework designed to enhance open-vocabulary semantic segmentation by leveraging multi-resolution VLM features. The exceptional quantitative and qualitative results obtained on well-established open-vocabulary semantic segmentation benchmarks serve as compelling evidence of its effectiveness and versatility. We hope our method can serve as a strong baseline for future research.
Appendix
A Overview
In the supplementary material for MROVSeg, we provide implementation detail (Sec. B), qualitative comparison (Sec. C), and limitation and discussion (Sec. D).
B Implementation Detail
B.1 Training Detail
During the training, we follow common practices [10, 36] to use AdamW optimizer with a base learning rate of $42e-$ 4 with poly decay schedule power $0.9$ to train the model, and the weight decay is set to $41e-$ 4. We use the same data augmentation strategies as in [36, 38, 22] for training, i.e., all training images firstly are randomly resized to [0.5,2.0]×[0.5,2.0$\times$[ 0.5 , 2.0 ] × of its original resolution, then are randomly cropped into the resolution of $640^{2}$. And for fair comparison, the test-time argumentation is not used.
B.2 Multi-grained Masked Attention
As introduced in Sec. 3.6, we enable the $\mathbf{X}_{\texttt{prop}}$ by perform masked cross attention with multi-resolution CLIP visual tokens. We show the pseudo code in Algorithm 1. During the inference, denote the number of CLIP token as $N$, the number of CLIP high-resolution token $4N$, the number of query token as $N_{\texttt{prop}}$, then the time complexity of the cross attention implementation is $\mathcal{O}(5N^{2}+20N_{\texttt{query}}N)$.
C Qualitative Results
In this section, we provide the segmentation results on COCO-Stuff val [3] (close-vocabulary setting), ADE [43] and Pascal Context [26] (open-vocabulary setting).
C.1 Comparison
We provide more qualitative comparison of our method MROVSeg with state-of-the-art methods SAN [36] and EBSeg [30] in Fig.S10. We observe that MROVSeg can handle complex scenarios while preserving spatial details (object boundaries) and extract global context effectively.
C.2 COCO
We visualize the closed-vocabulary semantic predictions of COCO-Stuff [3]. As shown in Fig.S9, our method is able to handle complex scenarios and provide precise mask predictions. However, as the labels exist semantically containment relationships, our model struggles to distinguish the regions with similar semantics. For instance, the model cannot find out the donut out of the food-other in fifth row of Fig.S9(b).
C.3 ADE
We visualize the open-vocabulary semantic segmentation predctions of MROVSeg on ADE [43] dataset with 150 and 847 categories benchmark respectively. As the result, we find MROVSeg can segment and identify the novel classes that only appear in 847-class annotations. As illustrated in Fig.S11, the model is able to segment and recognize phone booth and snow in (b) and (c) of Fig.S11 respectively.
Algorithm 1 Pseudocode of Multi-grained Masked Attention in Pytorch-like style.
1# visual_tokens: multi-resolution CLIP visual token with shape (5L, N, C)
2# M_local: local attention mask with shape [4L,N,C]
3# M_global: global attention mask with shape [L,N,C]
4# X_prop: [CLS] token with position embedding
5M = torch.cat([M_global, M_local], dim=0)
6# load frozen VLM layers
7for layer in cross_attn_layers:
8 X_prop = layer(X_prop,visual_tokens,attn_mask=M)
9visual_embedding = matmul(X_prop, W_proj)
10logits = bmm(visual_embedding, text_embedding.T())
C.4 Pascal Context
The qualitative results of segmentation performance of MROVSeg for 59 and 459 semantic categories on Pascal Context [26] dataset are shown in Fig.S12. Note that MROVSeg is able to segment small objects (such as the video camera in Fig.S12(c)), which is benefit from the high resolution VLM features and strong local contexts.
D Disucssion
In this section, we further discuss the limitation, future work and potential negtive impact of MROVSeg.
D.1 Limitation and Future Work
Due to computational constraints and for fair performance comparison, we conduct our experiments solely based on the CLIP [28] model. Other vision-language models, such as ALIGN [19], should also be relevant to this work.
This work involves tuning pretrained Vision-Language Models (VLMs) in a parameter-efficient manner, while pretrained VLMs often contain a large number of parameters, which significantly impacts the efficiency of segmentation. As mentioned in Sec.3 of the main paper, our method requires the use of a sliding window to repeatedly sample the same image multiple times during the inference, which, though affordable (demonstrated in Sec. 4.3), is also computationally unfriendly.
As shown in the Fig.S9(b), MROVSeg struggles to distinguish and process semantics with containment relationships, often leading to inaccurate recognition. For future work, introducing semantic priors based on category names is an promising direction to solve this issue. The datasets used in this work suffer from imprecise semantic mask annotations (as illustrated in Fig.S12(d), the cups on the table are not annotated), which may affect the real running performance of the model. Additionally, the evaluation method for open-vocabulary semantic segmentation has inherent ambiguities as shown in the Fig.S12(c). The ground mask is also reasonable to be classified to road as a truck drving on it. Recently, there have been some studies (such as  [1]) exploring the performance of open vocabulary semantic segmentation models across a wider range of categories and more specialized data. Thus, investigating more reasonable evaluation methods for open-vocabulary semantic segmentation is also a focus for future work.
D.2 Dataset usage
This work involves training segmentation models on public datasets(COCO [23], VOC [15], and ADE [43]). These datasets may contain personally identifiable information (such as facial characteristics), but they were responsibly collected and used. These datasets have benefited the image segmentation community for years.
D.3 Potential Negtive Impact
This work expands the scope of open-vocabulary semantic segmentation methods, offering significant advantages in areas like healthcare and autonomous driving. However, there is potential for misuse in applications such as surveillance, a common issue in most semantic segmentation research. Implementing regulatory measures to control the application of these algorithms might be an effective solution.
Figure S9: More Qualitative comparison with SAN [36] and EBSeg [30].
Figure S10: Qualitative results on COCO-Stuff [3] val. GT refer to groundtruth, Pred. refer to semantic prediction. We visualize the segmentation prediction of MROVSeg under complex real-world scenarios, such as muti-object scenes, low-light environment, and gray-scale photos.
Figure S11: Qualitative results on ADE [43] val. Pred/GT-150/847 refer to the prediction/groundtruth for 150/847 semantic categories. We visualize the semantic prediction of a image on 150 and 847 categories.
Figure S12: Qualitative results on Pascal Context [26] val. Similar to Fig.S11, we visualize the prediction and groundtruth for 59 and 459 semantic categories respectively.